A weight is added to a piston so that the volume of the gas inside the container is reduced from 3.0 L to 1.0 L at a constant temperature. How much heat do you need to add to the system if you want to increase the volume back to 3.0 L at this new pressure? Assume an ideal gas and an initial pressure of 1.0 atm
Answer #1

2? 34 4 Initial Pressure at latm Ideal Gas PV = P₂ V₂ = nRT. Samuvalue. P2 = x 3 atm. W12 = (201325) (Bc.) ln(43) J W2 = - 33P( VYD Now heat is applied Means Piston will move up Against const Pressure of Palm & Pweight = 3 atm. - W₂ = 3(3-1) (101325)AU=0 const Temperature. 812 = W 12 = -333.9515 Overall Prouss wanitaris 2 1 AU = EQ - EW ! AU=0 (As it Reforended to its InitThank you

Please let me know if you need anything else from me

Upvote if helpful

Have a great day